# Preference-driven Knowledge Distillation for Few-shot Node Classification

### v1

@@ E.1 Visualization
(a) Original student GNN
(b) GCNII
(c) EGNN
(d) LLMGNN
(e) GAugLLM
(f) AGST
(g) KDGA
(h) MSKD
(i) BGNN
(j) MTAAM
(k) FairGKD
(l) PKD # LN 5
@@ E.2 Ablation Study
The results on Cora and Amazon Ratings of our RL-based method (PKD) and the other three teacher selection methods (entropy-based ranking, i.e., selecting the teacher GNN with the highest prediction confidence, random selection, and end-to-end learning) are shown in Table 16. It is obvious that the RL-based method significantly outperforms the other three methods.
@@ E.3 Hyperparameters Sensitivity Analysis
As mentioned in Sec. 4.4, we perform the hyperparameter sensitivity analysis over the loss-weight coefficients $\alpha,\beta,\gamma,\eta$ on two datasets, and report the results in Tables 17 18 19 20. As result, the proposed PKD can achieve much better performance when $\alpha=0.5,\beta=1,\gamma=0.1,\eta=0.3$.
@@
$\alpha$ 2 1 0.9 0.8 0.7 0.6 0.5 0.4 0.3
Cora 70.12 69.49 72.71 69.05 72.41 73.15 75.66 74.70 73.78
Amazon Ratings 61.74 61.66 60.47 60.28 61.30 62.76 63.83 62.86 61.81
@@
$\beta$ 2 1 0.5 0.25
Cora 74.15 75.36 67.17 64.10
Amazon Ratings 63.93 63.91 63.23 62.94
@@
$\gamma$ 2 1 0.5 0.2 0.1 0.05
Cora 62.66 65.06 64.40 63.77 70.12 68.90
Amazon Ratings 63.93 63.93 63.63 63.71 64.01 62.58
@@
$\eta$ 0.9 0.7 0.5 0.3 0.1
Cora 89.42 86.67 88.43 90.27 89.96
Amazon Ratings 63.81 63.75 64.79 65.93 64.87

### v3

@@ E.2 Ablation Study
The results on Cora and Amazon Ratings using our PKDLlama (RL-based method) and the other three teacher selection methods (Entropy-based ranking, i.e., selecting the teacher GNN with the highest prediction confidence, Random selection, and End-to-end learning) are shown in Table 16. It is obvious that the RL-based method significantly outperforms the other three methods.
@@ E.3 Hyperparameters Sensitivity Analysis
As mentioned in Sec. 4.4, we perform the hyperparameter sensitivity analysis over the loss-weight coefficients $\alpha,\beta,\gamma,\eta$ on two datasets, and report the results in Tables 171819 and 20. As result, the proposed PKD can achieve much better performance when $\alpha=0.5,\beta=1,\gamma=0.1,\eta=0.3$.
@@
$\alpha$ 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1 2
Cora 73.78 74.70 75.66 73.15 72.41 69.05 72.71 69.49 70.12
Amazon Ratings 61.81 62.86 63.83 62.76 61.30 60.28 60.47 61.66 61.74
@@
$\beta$ 0.25 0.5 1 2
Cora 64.10 67.17 75.36 74.15
Amazon Ratings 62.94 63.23 63.91 63.93
@@
$\gamma$ 0.05 0.1 0.2 0.5 1 2
Cora 68.90 70.12 63.77 64.40 65.06 62.66
Amazon Ratings 62.58 64.01 63.71 63.63 63.93 63.93
@@
$\eta$ 0.1 0.3 0.5 0.7 0.9
Cora 89.96 90.27 88.43 86.67 89.42
Amazon Ratings 64.87 65.93 64.79 63.75 63.81